OpenStudy (anonymous):

Currently, Sam's coach is twice as old as Sam. When Sam started soccer 10 years ago, his coach was four times his age. How old is Sam now?

OpenStudy (anonymous):

You need to write two equation, can you think of what these equations could be?

OpenStudy (anonymous):

nope im dumb with math

OpenStudy (anonymous):

Let C=coach's age Let S=Sam's age write an expression using C and S for the first statement. Coach's age is twice as big as Sam's age.

OpenStudy (anonymous):

try!

OpenStudy (anonymous):

c(2)+s= ???? @juantweaver

OpenStudy (anonymous):

almost 2*age of Sam=age of Coach so 2S=C

OpenStudy (anonymous):

o alright

OpenStudy (anonymous):

Now for the second equation. Ten years ago Sam was what age?

OpenStudy (anonymous):

we are looking for a formula for this as well

OpenStudy (anonymous):

four times his age @juantweaver

OpenStudy (anonymous):

first we have to figure out Sam's age ten years ago This is S-10

OpenStudy (anonymous):

Then four times Sam's age ten years ago 4*(S-10)=(C-10)

OpenStudy (anonymous):

So the two equations are 4*(S-10)=C-10 C=2*S

OpenStudy (anonymous):

so what would i do next @juantweaver

OpenStudy (anonymous):

substitute second equation into the first. (Replace all C's in first equation with what C is equal to in the 2nd equation

OpenStudy (anonymous):

so its 4*(s-10)=2*s ? @juantweaver

OpenStudy (anonymous):

close 4*(S-10) = 2*S - 10

OpenStudy (anonymous):

alright can i solve it now ? or is there 1 more step ? @juantweaver

OpenStudy (anonymous):

just solve for S

OpenStudy (anonymous):

so its 4s-40=2*s-10

OpenStudy (anonymous):

yes, keep going

OpenStudy (anonymous):

4s-40=2s-20?

OpenStudy (anonymous):

@juantweaver

OpenStudy (anonymous):

no, return to 4s-40=2s-20 do you know how to combine like terms

OpenStudy (anonymous):

8s=800 ? id that right

OpenStudy (anonymous):

not quite almost though, instead of multiplying the numbers over, you should be adding them or subtracting them 4s-2s = -20+40

OpenStudy (anonymous):

o ok , then its 2s=60 right ?